# Deflickering Vision-Based Occupancy Networks through Lightweight Spatio-Temporal Correlation

### v3

@@ 7       Retrieve $s_{t-L:t-1}$, $m_{t-L:t-1}$, and $s_{t}$ from $\mathcal{W}$;
10       Decode attention outputs into correction results via Eq. (10);
11       Refine initial result prediction via Eq. (3);
12       $t\leftarrow t+1$;
@@
The resulting difference images $M_{t-L:t}$ are then used as input to the motion encoding module. Since these frame difference images typically share the same resolution as the original RGB images – often high-resolution [43] (e.g., $1600\times 900$) – direct processing can be expensive and unsuitable for real-time applications. To address this, we first downsample $M_{t-L:t}$ by a factor of five to reduce the spatial dimensions. Moreover, unlike RGB images that contain complex textures and color variations, frame difference images primarily capture motion boundaries. Therefore, we employ a lightweight three-layer convolutional neural network (CNN) followed by batch normalization (BN) to efficiently extract motion cues. This network is denoted by $\Phi_{\rm ME}$. Formally, the resulting motion features are computed as:
$m_{t-L:t-1}={\rm Concat}(\Phi_{\rm ME}(M_{t-k}:k\in[1,\cdots,L]).$ (7)
@@
The extracted feature $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ are then fed into a shared tokenization encoder that transforms them into compact token sets. Specifically, a $1{\times}1$ convolution is applied to first reduce channel dimensions, followed by a spatial unfolding operation that divides each image into $p{\times}p$ patches. Within each patch, pixel-wise features are averaged to form a $v$-dimensional vector. This process is applied to each camera view independently and results in the following tokenization operation $\Theta_{\rm TE}$:
$\Theta_{\rm TE}(I)=\overline{\mathcal{U}(p,p)(I\circledast\Theta_{1\times 1})},$ (8)
@@
The final output $E$ is a set of sparse tokens with shape $[|V|\cdot\lfloor\frac{h}{p}\rfloor\cdot\lfloor\frac{w}{p}\rfloor,1,d_{token}]$, where each token retains localized spatial or motion semantics. This tokenized format is especially well-suited for attention-based fusion in our multi-stream integration module, as discussed next.
III-D Spatio-temporal correlation strategy
To efficiently exploit historical static and motion cues while avoiding redundant computation, we adopt a spatio-temporal correlation strategy based on dual multi-head attention operations. Unlike previous approaches that directly fuse all temporal inputs indiscriminately [13, 5], our method disentangles the correlation modeling process into two independent streams: one for static context and another for motion cues. These correlations are computed in a shared compact latent space, improving both interpretability and efficiency.
@@ III-D Spatio-temporal correlation strategy
Given the sparse tokens $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ extracted in Section III-C, we use $s_{t}$ (the current frame token) as the reference context. The goal is to model how past appearance and motion features relate to the current frame through dual multi-head attention mechanisms (MHAM) [27].
  • Historical static attention: We use the tokenized historical static features $\Theta_{\rm TE}(s_{t-L:t-1})$ as queries ($Q_{\rm sta}$), while the tokenized current feature $\Theta_{\rm TE}(s_{t})$ serves as both keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). Each component is projected into a shared latent space via independent learnable transformations. This MHAM produces the output $f_{\rm sta}$, capturing long-range spatial appearance correlations, as illustrated in the first column of Fig. 4.
  • Historical motion attention: Similarly, the motion tokens $\Theta_{\rm TE}(m_{t-L:t-1})$ are used as queries ($Q_{\rm mot}$), and the same current feature tokens $\Theta_{\rm TE}(s_{t})$ are reused as keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). The resulting attention output $f_{\rm mot}$ captures temporally aligned motion dynamics, as shown in the second column of Fig. 4.
@@
$f_{\rm sta}$ $=\mathrm{Softmax}\left(\frac{W_{\rm sta}^{Q}Q_{\rm sta}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{grad}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$ (9)
$f_{\rm mot}$ $=\mathrm{Softmax}\left(\frac{W_{\rm mot}^{Q}Q_{\rm mot}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{grad}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$
where $W^{Q}_{*}$, $W^{K}_{*}$, and $W^{V}_{*}$ are independent learnable projection matrices for each stream. The $\mathrm{Softmax}(\cdot)$ function is used to normalize the attention scores, and the scaling factor $\sqrt{d_{grad}}$ helps stabilize gradients during training [27].
@@
Feature decoding and correction.
@@
III-E Optimization
As shown in Fig. 3, the static encoder $\Phi_{\rm SE}$ is directly inherited from well-established VON architectures [18, 12, 16]. In contrast, the motion encoder $\Phi_{\rm ME}$ is a lightweight, trainable module specifically designed to extract temporal dynamics and compensate for missing or incomplete voxel predictions caused by occlusions or visual ambiguity. In scenarios where the static encoder $\Phi_{\rm SE}$ produces suboptimal outputs, $\Phi_{\rm ME}$ serves as a complementary component by injecting motion-aware corrections. Notably, the motion encoder is unlikely to dominate or collapse unless the frozen static features are already sufficiently complete to support accurate occupancy predictions on their own.
@@ III-E Optimization
To ensure a fair comparison and retain consistency with prior work, we adopt the original loss functions used by each benchmark during training. For the SurroundOcc [12] benchmark, we employ the following composite loss function:
@@
where $\mathcal{L}_{\rm ce}$ is a voxel-wise cross-entropy loss for occupancy classification [44], $\mathcal{L}_{\rm sem}$ enforces semantic consistency across space [45], $\mathcal{L}_{\rm geo}$ encourages geometric coherence between neighboring voxels [46].
For the Occ3D [6] benchmark, following the convention in ViewFormer [16], we apply the following loss:
@@
where $\mathcal{L}_{\rm focal}$ is the focal loss [47] for handling class imbalance, $\mathcal{L}_{\rm ls}$ is the lovasz softmax loss [48] for improving boundary-level segmentation, $\mathcal{L}_{\rm l1}$ is an L1 regression loss (scaled by $\lambda$) to supervise motion-induced voxel correction.
@@ IV-A Experimental settings
We evaluate on two established occupancy benchmarks: (i) the SurroundOcc benchmark[12, 43], which generates occupancy labels by voxelizing aggregated multi-frame LiDAR in the ego frame, using manually annotated 3D bounding boxes to assign semantic categories; and (ii) the Occ3D benchmark[6], which provides 0.4m-resolution voxel-level labels with occlusion states through automated LiDAR aggregation and mesh reconstruction. Both benchmarks cover 1,050 driving scenes, each with up to 40 timestamped frames and six synchronized cameras (front, front-left/right, back, back-left/right) at 1600$\times$900 resolution. In our experiments, we enhance single-frame baselines [18, 12, 16] by aggregating features from $L$ historical keyframes. Additionally, unlabeled intermediate frames from the “sweeps” folder [43] are utilized as motion information for OccLinker.
@@
Both the SurroundOcc and Occ3D benchmarks adopt the same 3D occupancy resolution of $(H,W,D)=(200,200,16)$ and utilize the same number of surround-view cameras, with $|V|=6$. However, they differ in semantic granularity: SurroundOcc defines $C=17$ semantic classes, while Occ3D provides $C=18$ for finer-grained occupancy labeling. Following the respective conventions [12, 16] for static feature dimensions, we adopt a feature map resolution of $(h,w)=(116,200)$ for SurroundOcc and $(h,w)=(32,88)$ for Occ3D.
We employ the $\Phi_{\rm ME}$ module with a temporal sampling interval of $\Delta l\approx 0.1$ seconds to extract motion features from raw nuScenes images at a resolution of $900\times 1600$. Within the $\Phi_{\rm ME}$ module, input images are first downsampled by a factor of 5, resulting in $180\times 320$ feature maps to reduce GPU memory consumption. Motion inputs are computed using either frame difference operations [42] or optical flow analysis [49]. The token dimension and the gradient scaling factor are both set to $d_{token}=d_{grad}=32$, with the number of temporal patches set as $p=6$. For the convolutional layer $\Theta_{1\times 1}$, the kernel size, stride, padding, input channels, and output channels are set to $(1,1)$, $1$, 00, 512, and $d_{token}$, respectively. In contrast, for $\Theta_{3\times 3\times 3}$, these values are set to $(3,3,3)$, $1$, $1$, $3C$, and $C$, respectively.
All experiments were conducted using four NVIDIA L20 GPUs with a batch size of 1. We employed the AdamW optimizer on both the SurroundOcc [12] and Occ3D [6] benchmarks, setting the learning rate to $2\times 10^{-4}$, weight decay to $1\times 10^{-2}$, and betas to $(0.9,0.999)$. The complete training process took approximately 24 epochs, corresponding to around 1.5 GPU-days.
@@ IV-B Evaluation Metrics
GMO and GSO classification
To assess the temporal consistency and occupancy accuracy of moving and static objects, we adopt the object categorization scheme from Cam4DOcc [40], dividing objects into two broad classes: General Moving Objects (GMO) and General Static Objects (GSO). Objects that remain stationary relative to the ground are classified as GSO, whereas those that move relative to the ground belong to the GMO category. This classification scheme enables separate and quantitative analysis of moving and static objects within the scene. Detailed semantic categories for each class are listed in Table I.
@@
Occupancy accuracy metric
For a rigorous and comprehensive evaluation across benchmarks, we employ Intersection over Union (IoU) and Mean Intersection over Union (mIoU) metrics, which are standard in 3D semantic occupancy prediction tasks [50, 51, 52, 53]. The mIoU scores are computed separately over three groups: all classes, GMO classes, and GSO classes.
@@
Let $\sigma_{i,n}^{(x,y,z)}$ denote the semantic label of the $n$-th voxel at spatial coordinates $(x,y,z)$ in frame $i$. Define the indicator function $\delta(e_{1},e_{2})=\mathbb{I}(e_{1}\neq e_{2})$, which equals 1 if the two inputs differ and 0 otherwise.
Voxel changes between frames $i$ and $j$ fall into two categories: “Moving Object Change” (MOC) and “Static Object Change” (SOC), defined as follows:
@@
Table II: Definitions of MOC and SOC voxel changes. $N_{mc}$ and $N_{sc}$ denote the number of corresponding voxels.
@@
$\Delta_{m}(i,j)$ $=\dfrac{1}{N_{mc}}\sum\limits_{n=1}^{N_{mc}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right),$ (13)
$\Delta_{s}(i,j)$ $=\dfrac{1}{N_{sc}}\sum\limits_{n=1}^{N_{sc}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right).$
The overall temporal consistency scores for moving and static objects, $S_{m}$ and $S_{s}$, are then computed by aggregating disparities over sequential frames:
$S_{m/s}=1-\dfrac{1}{M-1}\sum\limits_{k=1}^{M-1}\Delta_{m/s}(k,k+1),$ (14)
where $M$ denotes the total number of frames in the scene. Final consistency scores $\overline{S_{m}}$ and $\overline{S_{s}}$ are averaged across all scenes. Higher values indicate smoother and more temporally consistent occupancy predictions.

### v4

@@ 5   Extract motion features from the high-frequency motion cues: $m_{t}\leftarrow\Phi_{\rm ME}(M_{t})$;
10   Decode the attention outputs into a correction component via Eq. (11);
 // Stage 3: Occupancy Refinement
11   Refine the initial occupancy prediction via Eq. (4);
12 Advance to the next timestep: $t\leftarrow t+1$;
@@
III-B Sparse tokenization for query construction
Static feature encoder
We extract static spatial features from both the current and previous keyframes by directly reusing the 2D vision backbone modules, denoted as $\Phi_{\rm SE}$, from foundational vision-based occupancy networks (e.g., SurroundOcc [surroundOcc] or ViewFormer [viewformer]). Let $S_{t}$ represent the current surround-view image and $S_{t-L:t-1}$ denotes the historical keyframes within a temporal window of size $L$. The corresponding static features are computed as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{t-L:t}=\Phi_{\rm SE}^{*}(S_{t-L:t})$}, (5)
where $\Phi_{\rm SE}^{*}$ shares parameters across all time steps.
Motion feature encoder
Although occupancy prediction is typically performed using keyframes sampled at a fixed rate, the ego-vehicle continuously captures high-frequency RGB images. These intermediate frames, located between consecutive keyframes, contain rich motion cues that are valuable for spatial-temporal perception. To efficiently extract motion information, we adopt a simple yet effective frame-difference technique [Frame_difference]. For each pair of adjacent keyframes $(S_{i},S_{i+1})$ within the historical window [tL,t][t-L,t], where $t-L\leq i<i+1\leq t$. We compute the difference between any two intermediate frames indexed by $(\tau_{a},\tau_{b})$ satisfying $i\leq\tau_{a}<\tau_{b}\leq i+1$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{i:i+1}=\{S^{\tau_{b}}_{i:i+1}-S^{\tau_{a}}_{i:i+1}\}$}. (6)
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M_{t-L:t-1}=\text{Concat}\left(\left\{F_{i:i+1}\mid i\in[t-L,t-1]\right\}\right)$}, (7)
where $\text{Concat}(\cdot)$ denotes the concatenation operation performed along the temporal dimension, combining motion cues from all intervals in the historical window.
The resulting difference images $M_{t-L:t}$ are then used as input to the motion encoding module. Since these frame difference images typically share the same resolution as the original RGB images – often high-resolution [nuScenes] (e.g., $1600\times 900$) – direct processing can be expensive and unsuitable for real-time applications. To address this, we first downsample $M_{t-L:t-1}$ by a factor of five to reduce the spatial dimensions. Moreover, unlike RGB images that contain complex textures and color variations, frame difference images primarily capture motion boundaries. Therefore, we employ a lightweight three-layer convolutional neural network (CNN) followed by batch normalization (BN) to efficiently extract motion cues. This network is denoted by $\Phi_{\rm ME}$. Formally, the resulting motion features are computed as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m_{t-L:t-1}=\Phi_{\rm ME}(M_{t-L:t-1}).$} (8)
@@ III-B Sparse tokenization for query construction
The extracted feature $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ are then fed into a shared tokenization encoder that transforms them into compact token sets. Specifically, a $1{\times}1$ convolution is applied to first reduce channel dimensions, followed by a spatial unfolding operation that divides each image into $p{\times}p$ patches. Within each patch, pixel-wise features are averaged to form a $v$-dimensional vector. This process is applied to each camera view independently and results in the following tokenization operation $\phi$:
$\phi(I)=\overline{\mathcal{U}(p,p)(I\circledast\Theta_{1\times 1})},$ (9)
@@
The final output $E$ is a set of sparse tokens with shape $|V|\cdot\lfloor\frac{h}{p}\rfloor\cdot\lfloor\frac{w}{p}\rfloor,1,d_{token}]$, where each token retains localized spatial or motion semantics. This tokenized format is especially well-suited for attention-based fusion in our multi-stream integration module, as discussed next.
III-C Spatio-temporal correlation strategy
To efficiently exploit historical static and motion cues while avoiding redundant computation, we adopt a spatio-temporal correlation strategy based on dual multi-head attention operations. Unlike previous approaches that directly fuse all temporal inputs indiscriminately [bevdet4d, yang2024adaptiveocc], our method disentangles the correlation modeling process into two independent streams: one for static context and another for motion cues. These correlations are computed in a shared compact latent space, improving both interpretability and efficiency.
@@ III-C Spatio-temporal correlation strategy
Given the sparse tokens $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ extracted in Section III-B, we use $s_{t}$ (the current frame token) as the reference context. The goal is to model how past appearance and motion features relate to the current frame through dual multi-head attention mechanisms (MHAM) [attention_is_all_you_need].
  • Historical static attention: We use the tokenized historical static features $\phi(s_{t-L:t-1})$ as queries ($Q_{\rm sta}$), while the tokenized current feature $\phi(s_{t})$ serves as both keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). Each component is projected into a shared latent space via independent learnable transformations. This MHAM produces the output $f_{\rm sta}$, capturing long-range spatial appearance correlations, as illustrated in the first column of Fig. 4.
  • Historical motion attention: Similarly, the motion tokens $\phi(m_{t-L:t-1})$ are used as queries ($Q_{\rm mot}$), and the same current feature tokens $\phi(s_{t})$ are reused as keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). The resulting attention output $f_{\rm mot}$ captures temporally aligned motion dynamics, as shown in the second column of Fig. 4.
@@
$f_{\rm sta}$ $=\mathrm{Softmax}\left(\frac{W_{\rm sta}^{Q}Q_{\rm sta}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$ (10)
$f_{\rm mot}$ $=\mathrm{Softmax}\left(\frac{W_{\rm mot}^{Q}Q_{\rm mot}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$
where $W^{Q}_{*}$, $W^{K}_{*}$, and $W^{V}_{*}$ are independent learnable projection matrices for each stream. The $\mathrm{Softmax}(\cdot)$ function is used to normalize the attention scores, and the scaling factor $\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}$ helps stabilize gradients during training [attention_is_all_you_need].
@@
Feature decoding and correction
@@
Notably, we do not rely on any explicit coordinate transformation between frame $t$ and historical frames tL:t1t-L:t-1. Unlike many occupancy prediction methods [xu2025exploiting, viewformer] that incorporate ego-motion information (e.g., relative translation and rotation from IMU or odometry) for geometric alignment or implicit motion compensation, OccLinker directly associates historical static and dynamic cues in the sparse latent space. This design keeps OccLinker lightweight, sensor-agnostic, and fully compatible with both single-frame and history-aware VONs, without requiring additional motion supervision or modifying their original coordinate systems.
III-D Optimization
As shown in Fig. 3, the static encoder $\Phi_{\rm SE}$ is directly inherited from well-established VON architectures [MonoScene, surroundOcc, viewformer]. In contrast, the motion encoder $\Phi_{\rm ME}$ is a lightweight, trainable module specifically designed to extract temporal dynamics and compensate for missing or incomplete voxel predictions caused by occlusions or visual ambiguity. In scenarios where the static encoder $\Phi_{\rm SE}$ produces suboptimal outputs, $\Phi_{\rm ME}$ serves as a complementary component by injecting motion-aware corrections. Notably, the motion encoder is unlikely to dominate or collapse unless the frozen static features are already sufficiently complete to support accurate occupancy predictions on their own.
@@ III-D Optimization
To ensure a fair comparison and retain consistency with prior work, we adopt the original loss functions used by each benchmark during training. For the SurroundOcc [surroundOcc] benchmark, we employ the following composite loss function:
@@
where $\mathcal{L}_{\rm ce}$ is a voxel-wise cross-entropy loss for occupancy classification [cross_entropy_loss_2], $\mathcal{L}_{\rm sem}$ enforces semantic consistency across space [Semantic_loss], $\mathcal{L}_{\rm geo}$ encourages geometric coherence between neighboring voxels [geometry_affinity_loss].
For the Occ3D [Occ3D] benchmark, following the convention in ViewFormer [viewformer], we apply the following loss:
@@
where $\mathcal{L}_{\rm focal}$ is the focal loss [focal_loss] for handling class imbalance, $\mathcal{L}_{\rm ls}$ is the lovasz softmax loss [lovasz_loss] for improving boundary-level segmentation, $\mathcal{L}_{\rm l1}$ is an L1 regression loss (scaled by $\lambda$) to supervise motion-induced voxel correction.
@@ IV-A Experimental settings
[…]
@@
Both the SurroundOcc and Occ3D benchmarks adopt the same 3D occupancy resolution of $(H,W,D)=(200,200,16)$ and utilize the same number of surround-view cameras, with $|V|=6$. However, they differ in semantic granularity: SurroundOcc defines $C=17$ semantic classes, while Occ3D provides $C=18$ for finer-grained occupancy labeling. Following the respective conventions [surroundOcc, viewformer] for static feature dimensions, we adopt a feature map resolution of $(h,w)=(116,200)$ for SurroundOcc and $(h,w)=(32,88)$ for Occ3D. Notably, both $s_{t-L:t}$ and $m_{t-L:t}$ share the same dimensionality of $|V|\times c\times w\times h$.
We employ the $\Phi_{\rm ME}$ module with a temporal sampling interval of $\Delta l\approx 0.1$ seconds to extract motion features from raw nuScenes images at a resolution of $900\times 1600$. Within the $\Phi_{\rm ME}$ module, input images are first downsampled by a factor of 5, resulting in $180\times 320$ feature maps to reduce GPU memory consumption. Motion inputs are computed using either frame difference operations [Frame_difference] or optical flow analysis [OpticalFlow]. The token dimension and the gradient scaling factor are both set to $d_{token}=d_{grad}=32$, with the number of temporal patches set as $p=6$. For the convolutional layer $\Theta_{1\times 1}$, the kernel size, stride, padding, input channels, and output channels are set to $(1,1)$, 11, 0, 512, and $d_{token}$, respectively. In contrast, for $\Theta_{3\times 3\times 3}$, these values are set to $(3,3,3)$, 11, 11, $3C$, and $C$, respectively. OccLinker is integrated on top of this baseline to complement the missing inter-frame motion cues, rather than replacing any component of the original design.
All experiments were conducted using four NVIDIA L20 GPUs with a batch size of 1. We employed the AdamW optimizer on both the SurroundOcc [surroundOcc] and Occ3D [Occ3D] benchmarks, setting the learning rate to $2\times 10^{-4}$, weight decay to $1\times 10^{-2}$, and betas to $(0.9,0.999)$. The overall training took approximately 24 epochs, corresponding to about 1.5 GPU-days in total, including around 1.4 GPU-days for pretraining the base VON and 0.1 GPU-day for fine-tuning OccLinker.
@@ IV-B Evaluation Metrics
Occupancy accuracy metric
For a rigorous and comprehensive evaluation across benchmarks, we employ Intersection over Union (IoU) and Mean Intersection over Union (mIoU) metrics, which are standard in 3D semantic occupancy prediction tasks [PASCAL, Microsoft_COCO, Cityscapes_dataset, Mask_R_CNN]. IoU measures the global overlap between all predicted occupied voxels and their ground-truth counterparts, regardless of semantic categories. It reflects the overall spatial accuracy of the 3D occupancy prediction. In contrast, mIoU is the mean IoU computed by averaging IoUs over all semantic classes, which reflects class-balanced semantic accuracy.
To further evaluate the temporal consistency and occupancy accuracy for both moving and static objects, we adopt the object categorization scheme from Cam4DOcc [Cam4docc], dividing objects into two broad classes: General Moving Objects (GMO) and General Static Objects (GSO). Objects that remain stationary relative to the ground are classified as GSO, whereas those that move relative to the ground belong to the GMO category. This classification scheme enables a fine-grained and quantitative analysis of moving and static objects within the scene. Detailed semantic categories for each class are listed in Table I. The mIoU scores are reported separately for three groups: all classes, GMO classes, and GSO classes.
@@
Let $\sigma_{i,n}^{(x,y,z)}$ denote the semantic label of the $n$-th voxel at spatial coordinates $(x,y,z)$ in frame $i$. Define the indicator function $\delta(e_{1},e_{2})=\mathbb{I}(e_{1}\neq e_{2})$, which equals 1 if the two inputs differ and 0 otherwise. Voxel changes between frames $i$ and $j$ fall into two categories: “Moving Object Change” (MOC) and “Static Object Change” (SOC), defined as follows:
@@
Table II: Definitions of MOC and SOC voxel changes.
@@
$\Delta_{m}(i,j)$ $=\dfrac{1}{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}}\sum\limits_{n=1}^{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right),$ (14)
$\Delta_{s}(i,j)$ $=\dfrac{1}{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}}\sum\limits_{n=1}^{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right).$
where $N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}$ and $N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}$ represent the number of corresponding voxels in the moving and static regions, respectively.
The overall temporal consistency scores for moving and static objects, $S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}$ and $S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}$, are then computed by aggregating disparities over sequential frames:
$S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m/s}}=1-\dfrac{1}{M-1}\sum\limits_{k=1}^{M-1}\Delta_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m/s}}(k,k+1),$ (15)
where $M$ denotes the total number of frames in the scene. Final consistency scores $\overline{S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}}$ and $\overline{S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}}$ are averaged across all scenes. Higher values indicate smoother and more temporally consistent occupancy predictions.